The following equations describe an economy (think of C, I, G, etc as being measured in billions and i as a percentage; a 5 percent interest rate implies i = 5)

           

C = 0.8 (1 – t) Y

           t = 0.25

           I = 900 – 50i

           G = 800

           L = 0.25Y – 62.5.i

           M / P = 500


Show that a given change in the money stock has a larger effect on output the less interest sensitive is the demand for money.



1
Expert's answer
2021-07-04T17:27:46-0400

If the demand for money is extremely sensitive to interest rates, then the rate of increase in interest rates is small, the demand for money will decrease, and the money market will return to equilibrium.

Y=C+I+G=0.8(1-t)Y+900-50R+800=0.6Y+900-50R+800=0.6Y+1700.8-50R

Y- 0.6Y=1700-50R

0.4Y =1700-50R

Y=4250-125R

0.25Y-62.5R=500

0.25Y-500=62.5R

R=0.004Y-8

Y=4250-125(0.004Y-8)

Y=4250-0.5Y+1000

Y+0.5Y=5250

1.5Y=5250

Y=3500

R=0.004×3500-8=6